OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? 6y = 9x2

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

Please help

OpenStudy (dtan5457):

They don't give you a x and y value?

OpenStudy (anonymous):

no

OpenStudy (dtan5457):

Not sure to be honest.. only thing I can think of is 6y=9x^2 divide by 6 y=3/2x^2 constant (k)=3/2?

OpenStudy (anonymous):

Thats what i got thank you!

OpenStudy (dtan5457):

Is that an answer choice?